OpenStudy (anonymous):

check plz!!!!! Simplify completely: 5x^3+10x^2+15x/5x

OpenStudy (anonymous):

answer: x^4 + 2x^3 + 3x^2

OpenStudy (anonymous):

You just need to divide everything on 5x: 5x^3/5x=x^2 10x^2/5x=2x 15x/5x=3 Your answer is: x^2+2x+3

OpenStudy (anonymous):

can you check another for me?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Simplify completely: 8x^3+6x^2-4x/-2x

OpenStudy (anonymous):

answer: −4x^2 − 3x + 2

OpenStudy (anonymous):

8x^3/2x + 6x^2/2x - 4x/2x -4^2 - 3x + 2 Hope this Helps!

OpenStudy (anonymous):

it does thnx :)

OpenStudy (anonymous):

np ally just message me for help oki i love to help

OpenStudy (anonymous):

allie*

OpenStudy (anonymous):

okay :) i have 1 last one but ill message it to u :)

OpenStudy (anonymous):

are you doing a 5 question test

OpenStudy (anonymous):

yea lol do u have 5?

OpenStudy (anonymous):

u do connections academy too?

OpenStudy (anonymous):

lol yea i sis that test

OpenStudy (anonymous):

is this the test 1. Simplify completely: 12x^3-4x^2+8x over -2x a. -6x^2+2x+4 b. -6x^2+2x-4 c. 6x^4-2x^3+4x^2 d. 6x^2-4x^2+8x 2.Simplify completely: 3x^4+5x^2+2x over x a. 3x^3+5x+2x over x b. 3x^3+5x+2 c. 3x^2+5x^2+2x d. 3x^5+5x^3+2x^2 3. Simplify completely: 5x^3+10x^2+15x over 5x a. x^2+10x^2+15x b. 2x+6 c. x^2+2x+3 d. x^4+2x^3+3x^2 4. Simplify completely: 14x^5y^4+21x^3y^2 over 7x^3y a. 2x^8y^5+3x^6y^3 b. 2x^2y^4+3xy c. 2x^2y^3+3y d. 2x^2y^3+3y^2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

do u know the answer to this question? Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

yea i dot stuck on that one

OpenStudy (anonymous):

got*

OpenStudy (anonymous):

okay ill post it and see if i can get someelse to answer thnx though :)

OpenStudy (anonymous):

np